OpenStudy (anonymous):

Polynomials and Polynomial Functions Unit Test lesson 11 unit 5

OpenStudy (anonymous):

1. Which equation is most likely represented by the graph above? (1 point) Correct answer (1 pt) y = (2x – 1)(2x – 3)(2x + 5) Unselected answer (0 pts) y = (2x + 1)(2x + 3)(2x – 5) Unselected answer (0 pts) y = –(2x + 1)(2x + 3)(2x – 5) Unselected answer (0 pts) y = (2x + 1)(2x + 3)(2x + 5) 1 /1 point 2. Shown below is the graph of . What are the apparent zeros of the function graphed above? (1 point) Unselected answer (0 pts) {–0.7, –2.7} Unselected answer (0 pts) {–10.2, 10.2} Unselected answer (1 pt) {–2, 1, 4} Incorrect Answer (0 pts) {–4, –1, 2} 0 /1 point 3. What is a polynomial function in standard form with zeros 1, 2, –2, and –3? (1 point) Unselected answer (0 pts) x4 + 2x3 + 7x2 – 8x + 12 Correct answer (1 pt) x4 + 2x3 – 7x2 – 8x + 12 Unselected answer (0 pts) x4 + 2x3 – 7x2 + 8x + 12 Unselected answer (0 pts) x4 + 2x3 + 7x2 + 8x + 12 1 /1 point Multiple Choice 4. What are the zeroes of the function? What are their multiplicities? f (x) = 4x3 – 20x2 + 24x (1 point) Incorrect Answer (0 pts) The numbers –3, –2, and 0 are zeroes of multiplicity 1. Unselected answer (0 pts) The numbers 3, 2, and 0 are zeroes of multiplicity 2. Unselected answer (0 pts) The numbers –3, –2, and 0 are zeroes of multiplicity 2. Unselected answer (1 pt) The numbers 3, 2, and 0 are zeroes of multiplicity 1. 0 /1 point 5. What are the real or imaginary solutions of the polynomial equation?x3 – 64 = 0 (1 point) Unselected answer (0 pts) 4, –1 + , –1 + Incorrect Answer (0 pts) 4, 1 + , 1 + Unselected answer (1 pt) 4, –2 + 2i3, –2 – 2i3 Unselected answer (0 pts) 4, –2 + , 2 + 0 /1 point 6. Which is equivalent to the following expression? (1 point) Incorrect Answer (0 pts) 3p2 + 8pq – 3q2 Unselected answer (0 pts) 4p2 + 8pq + q2 Unselected answer (0 pts) 4p4 + 8pq – 3q4 Unselected answer (1 pt) 4p2 + 8pq – 3q2 0 /1 point 7. Which is a factored form of 8x3 + 27? (1 point) Unselected answer (0 pts) (2x + 3)(2x + 3)(2x + 3) Unselected answer (1 pt) (2x + 3)(4x2 – 6x + 9) Incorrect Answer (0 pts) (2x – 3)(4x2 + 6x + 9) Unselected answer (0 pts) (2x – 9)(4x2 + 18x +81) 0 /1 point Multiple Choice 8. Divide –3x3 – 4x2 + 4x +3 by x – 2. (1 point) Incorrect Answer (0 pts) –3x2 + 2x + 24 Unselected answer (1 pt) –3x2 – 10x – 16, R – 29 Unselected answer (0 pts) –3x2 – 10x – 16 Unselected answer (0 pts) –3x2 + 2x + 24, R 35 0 /1 point 9. Find the roots of the polynomial equation. x3 – 4x2 + x + 26 (1 point) Unselected answer (1 pt) 3 + 2i, 3 – 2i, –2 Unselected answer (0 pts) 3 + 2i, 3 – 2i, 2 Unselected answer (0 pts) –3 + 2i, –3 – 2i, 2 Incorrect Answer (0 pts) –3 + 2i, –3 – 2i, –2 0 /1 point 10. Which correctly describes the roots of the following cubic equation? x3 – 5x2 + 3x + 9 = 0 (1 point) Unselected answer (0 pts) Three real roots, each with a different value Unselected answer (0 pts) One real root and two complex roots Unselected answer (1 pt) Three real roots, two of which are equal in value Incorrect Answer (0 pts) Two real roots and one complex root 0 /1 point 11. What is the second term of (s + v)5? (1 point) Unselected answer (0 pts) 25s4 Incorrect Answer (0 pts) 25s4v Unselected answer (1 pt) 5s4v Unselected answer (0 pts) 125s4v 0 /1 point 12. One zero of f(x) = x3 – 2x2 – 5x + 6 is 1. What are other zeros of the function? (1 point) Unselected answer (1 pt) 3 and –2 Incorrect Answer (0 pts) –3 and 2 Unselected answer (0 pts) 1 and 2 Unselected answer (0 pts) 1 and –3 0 /1 point 13. Which of the following quartic functions has x = –2 and x = –3 as its only two real zeros? (1 point) Unselected answer (0 pts) x4 – 5x3 + 7x2 – 5x + 6 Correct answer (1 pt) x4 + 5x3 + 7x2 + 5x + 6 Unselected answer (0 pts) x4 + 5x3 + 7x2 – 5x – 6 Unselected answer (0 pts) x4 – 5x3 + 7x2 + 5x – 6